OpenStudy (taylovesmatthew):

Find the percent of increase from 360 to 530. Round to the nearest tenth of a percent if necessary.

OpenStudy (triciaal):

change is the difference and percent means per 100 530-360 = 170 170/360 = x/100 solve for x to 2 decimal places and then round to one decimal place